Calculate the number of BTU evolved when 50 lb of H2SO4 are added with 50 lb of H2O at 25 °C?
1
Expert's answer
2016-04-25T12:15:05-0400
Btu – British thermal unit
1 Btu = 1055.06 J = 1.06 kJ
1 lb = 0.454 kg
50 lb = 22.7 kg
M (H2SO4) = 98 g/mol
n (H2SO4) = 22700/98 = 232 mol
n (H2O) = 22700/18 = 1261 mol
Enthalpy of dissolution of H2SO4 at this temperature and at this proportion to water is -40 kJ/mol.
Therefore the amount of heat emitted is: 40*232 = 9080 kJ =8566 Btu